OpenStudy (anonymous):

The weight of water is 62 1/2 lb per cubic foot. What is the weight of 4 1/9 cubic feet of water?

OpenStudy (danjs):

use ratios of 1....

OpenStudy (anonymous):

How do I do that

OpenStudy (danjs):

\[\frac{ 37/9 ft^3 }{ 1 }\frac{ 62.5 lbs }{ ft^3 } = \]

OpenStudy (danjs):

see the ft^3 cancels and you are left with lbs of water

OpenStudy (danjs):

4 1/9 = 37/9

OpenStudy (anonymous):

ok

OpenStudy (danjs):

so multiply 37/9 by 62.5 to get ## lbs of water

OpenStudy (anonymous):

I thought that was right but wasn't sure if it was the right answer

OpenStudy (danjs):

yes, that is why you keep track of the units, the numbers are less important You start with 4 1/9 f^3 and you want pounds

OpenStudy (danjs):

\[\frac{ ft^3 }{ 1 }*\frac{ lbs }{ ft^3 } = lbs\]

OpenStudy (anonymous):

So the answer comes out to a decimal

OpenStudy (danjs):

62 1/2 * 4 1/9 = 125/2 * 37/9 = 4625/18

OpenStudy (danjs):

or a decimal with a calculator = about 257 lbs of water

OpenStudy (anonymous):

256.944444

OpenStudy (anonymous):

Thanks for your help

OpenStudy (danjs):

your welcome, anytime